# This Version: March 9, 2024

### v1

@@ 5 Analysis of Examples
In this section, we show the applicability of our results to Examples 2 and 3, where learning and decision-making are interlinked. Example 2 has an explicit analytical solution and is more of an illustrative example of the structure employed in this paper. However, Example 3, like most problems in dynamic programming, does not have a closed-form solution and, therefore, is most amenable to analysis with our framework.
@@
We now verify the prediction of our results vis-á-vis the analytical solution. A fall in the cost of effort $c$ is a positive shock, and therefore, leads to a higher inference of ability, $\theta^{*}=1-c,$ in the Berk-Nash equilibrium, as predicted in Theorem 2. Similarly, a fall in the cost of the effort leads to an increase in the probability of success at the steady state $m_{s}^{*}(1);$ since an increase in the model parameter is a positive shock. Here, the stationary distribution $m^{*}_{s}$ is increasing in the usual (first) stochastic order dominance in the cost of effort.
@@ 6 Welfare Comparisons
We now turn towards comparing the objective welfare of an agent facing a regular SMDP, under correctly specified and misspecified Bayesian learning and characterize an upper bound on the difference between the two instances, the costs of misspecification, in terms of the primitives of the environment. For the purposes of this section, we shall assume that the utility function is bounded and strictly concave. Further, we shall assume that $u:\mathbb{S}\times\mathbb{X}\rightarrow\mathbb{R}$ is continuously differentiable in actions.
Let $g(s,\bar{\theta}),$ parameterized by a parameter $\bar{\theta},$ denote an optimal policy function. Suppose it exists under both correctly specified and misspecified environments, where $\bar{\theta}$ take values $\theta^{*}$ and $\theta_{*}$ under correctly specified and misspecified learning, respectively. We are interested in comparing the welfare of a correctly specified agent with parameter $\theta^{*}$ to the welfare of an agent who settles at the parameter $\theta_{*}$ asymptotically, under misspecified Bayesian learning as in Equation (1). Then, the agent’s welfare under the parameter $\bar{\theta},$ is the objective ex-ante expected discounted payoff, $W(s,\bar{\theta})$ of choosing the optimal action, $g(s,\bar{\theta}),$ and is given as,
@@
We interpret the difference $\gamma$ between the two policy functions in the sup-norm as the approximation error in the space of policy functions. Then, for a given approximation error $\gamma,$ Equation (15) supplies an upper bound on the welfare comparison in terms of the primitives, namely the discount factor, and the absolute bounds on the utility, the marginal utility function, and the KL divergence. Notice that if $\gamma=0,$ that is, the policy functions under the two instances of learning are identical, then the two welfare quantities are equal, by construction. The upper bound in Equation (15) gives intuitive comparative statics, vis-a-vis the model primitives. For instance, the discrepancy in welfare is larger, the greater the approximation error, $\gamma.$ Similarly, for a positive approximation error, the discrepancy increases in the discount factor. That is, as the agent gets increasingly patient, reflected in a higher discount factor $\beta,$ his approximation error keeps accumulating over the horizon, leading to a greater discrepancy in welfare. Furthermore, higher bounds on the utility function and its corresponding marginal utility function lead to a greater discrepancy in the welfare. The marginal utility function plays a role in the discrepancy by attaching itself to the approximation error. Intuitively, getting the policy wrong is more pronounced if it matters at the margins. Notice that for the upper bound on the KL divergence, we only compare the implied distributions by the true transition function $Q,$ under the two policy functions. This is because our interest lies in comparing the objective welfare under $Q$.
It is worth mentioning here that bound in Equation (15) is not specific to a misspecified MDP, rather, it holds for any MDP environment with a endogenous state evolution process, when plugged with alternate policy functions. In this sense, it is related to Lemma 3.1 in Santos (2000) which does a similar exercise in bounding welfare discrepancies. However, the state process in there is exogeneous. Furthermore, Santos (2000) uses Euler residuals to provide upper bounds on the approximation error, in terms of the primitives, for both the policy and the value functions, which is useful in the context of computing numerical approximations. Often, MDP environments do not have closed form solutions, thus making economists rely on numerical techniques to solve for them computationally. We hope that Theorem 5 could be instructive in this regard, and help in the numerical approximation of the Berk-Nash equilibria.
@@ 7 Concluding Remarks and Extensions
Models, by their very nature, offer simplified abstractions of reality, inevitably omitting certain nuances. This paper delineates conditions under which an important qualitative property, the comparative statics of decision-making and the corresponding inference, is preserved under model misspecification. The main contribution of this paper lies in establishing monotone comparative statics results for misspecified dynamic optimization problems, by a novel application of techniques in the fixed points literature (Smithson (1971)) and used in the context of large dynamic economies (Acemoglu and Jensen (2015)). We also illustrate the utility of these results for general interest environments. Further, we provide an upper bound on the welfare between correct and misspecified learning, in terms of the primitives. We conclude by outlining several promising directions for extending our findings to more broader environments.
@@
Multi-dimensional parameter spaces. Since most of the current applications in the literature predominantly involve one-dimensional models, [^47] This feature of the current literature is also noted in Esponda, Pouzo, and Yamamoto (2021). the results presented in the paper assume the model (parameter) set to be a compact subset of the real line. However, they could be extended to multi-dimensional parameter spaces. This would involve a suitable rehabilitation of Assumptions 3 and 4 for lattice and non-lattice multi-dimensional parameter spaces. Although, one limitation of such a result using the current techniques is that the comparative statics results imply that all inferred parameters in the equilibrium must be monotonic together, which may not be ideal for specific economic applications.
@@
Other forms of updating. While our results primarily address the asymptotic behavior of Bayesian learners, it’s important to note that other inference processes, such as Maximum Likelihood Estimation (MLE) and moment-based learning, can lead to the same asymptotic inferences. Csaba and Szoke (2023) provide an illustrative example that demonstrates the identical nature of inference between Bayesian inference and methods based on likelihoods.
@@
Static environments and dynamic concerns. Our results are tailored for dynamic MDP environments but they potentially are also applicable, subject to modifications, to static environments, such as the one in Esponda and Pouzo (2016) for static games. This also spills over to settings where agents are endogenously concerned about their misspecified models, such as those in Lanzani (2022). [^48] The reader is referred to Corollary 1 in Lanzani (2022).
@@
Monotone comparative dynamics. The scope of our paper is limited to analyzing the comparative statics properties of the equilibrium objects in the steady state. An interesting question arises regarding how the dynamics of the misspecified learning process might react to changes in its primitives. Specifically, we do not know how the variations in the frequency of state-action pairs, as well as the sequence of posteriors as described in Equation (1), would be influenced because of changes in the economic primitives. There is added complexity in the dynamics due to potential complementarities between current actions and the inference process via the role of experimentation. In this regard, the tools developed in Balbus, Dziewulski, Reffett, and Woźny (2022) for monotone comparative dynamics for stochastic games could be useful for future exploration.
@@
Misspecification in large economies. Our results are also applicable to misspecified dynamic economies with a continuum of agents, such as the one conceived in Molavi (2019). Given that the techniques we rely on (Acemoglu and Jensen (2015)) were developed in the context of economies with a continuum of agents, the results in this paper should potentially be straightforward to apply on equilibrium concepts, Molavi (2019) develops for the boundedly rational agents working with misspecified models in macroeconomic environments.
@@ Appendix B Proofs
The proofs for Theorems 1 and 2 are structured in three steps, with the proof technique similar to that in Acemoglu and Jensen (2015). For Theorem 2, the first step involves showing that for any fixed model distribution $\mu\in\mathcal{M}_{1}(\Theta)$, the set of stationary distributions on states and actions, $m^{*}$, induced by the optimal policy correspondence $G,$ will be Type I (Type II) monotonic in the primitives $p$. The second step involves constructing a mapping $\hat{\theta}$ that for each given $\mu$ and $p$ yields a set of model distributions, $\mu^{\prime}s$. It is the fixed points of this map that are the equilibrium model distributions $\mu^{*},$ given $p$. Finally, the third step involves the least and greatest selections from this map will be increasing in $p.$ This in turn leads us to give a new existence proof of Theorem 1 that relies on the monotonicity and identification properties of the equilibrium map, $T$. The rest of the proofs (Theorems 3 and 4) shall follow an analogous structure. Proof of Theorem 5 is based on the Taylor expansion of the utility function and on an application of new
The equilibirum mapping $T$ associated with the Berk-Nash equilibrium (Definition 5) is a set-valued function on the product space of probability measure on states and actions, and parameter space, $T:W\rightarrow 2^{W},$ where $W=\mathcal{M}_{1}(\mathbb{S}\times\mathbb{X})\times\mathcal{M}_{1}(\Theta)$ such that $\mathcal{T}(m,\mu)=\mathcal{M}(m,\mu)\times\mathcal{M}\left(\Theta_{Q}(m)\right)$, where
@@
for any $\mu\in\mathcal{M}_{1}(\Theta),\mathcal{F}(\mu)$ is the set of all $m^{\prime}$ that satisfies the condition of optimality and stationarity as in Definition 5.
@@
Let the regular SMDP satisfy assumptions 1 and 2. Then for any given $p,$ the optimal policy correspondence $G:\mathbb{S}\times P\rightarrow 2^{\mathbb{X}}$ is increasing in the state $s$ in the strong set order. In particular, it has a least and a greatest selection, and they are increasing in the state $s$ [^50] We suppress the dependence on $\mathcal{M}_{1}(\Theta)$ for notational convenience.
@@
$G(s,\mu,p)=\operatornamewithlimits{argmax}\limits_{x\in\mathbb{X}}\displaystyle\ u(s,x)+\delta\int_{S}V(s^{\prime},\mu,p)\bar{Q_{\mu}}(\mathrm{d}s^{\prime}|s,x),$ (17)
where $\bar{Q}_{\mu}=\displaystyle\int_{\Theta}Q_{\theta}\mu(d\theta).$ Given Assumptions 1 and 2, and by the application of Theorem 3.9.2 in Topkis (1998), the expression on the right-hand side is supermodular in $(s,x)$. Therefore, the optimal policy correspondence will be increasing in $s$ in the strong-set order. Further, given the optimal correspondence, the greatest and least selection exist and are increasing in the state.
@@
We only prove the above statement for the greatest selection in the Type I case. [^51] The rest of the cases follow analogously. Consider probability measures, $\nu_{1},\nu_{2}\in\mathcal{M}(\mathbb{S}\times\mathbb{X})$ such that $\nu_{2}\succsim_{st}\nu_{1}.$ To prove that the fixed point correspondence $T$ is Type I monotone, we need to show that for any $\lambda_{1}\in$ $T\nu_{1}$, there exists a $\lambda_{2}\in T\nu_{2}$ such that $\lambda_{2}\succsim_{st}\lambda_{1}.$ That is, if $\lambda_{1}\in T\nu_{1},$ then there exists a measurable selection $g_{1}:\mathbb{S}\rightarrow\mathbb{X},$ such that, for all $A\times B\in\mathcal{B}(\mathbb{S}\times\mathbb{X})$,
@@
The first inequality follows from $\nu_{2}\succsim_{st}\nu_{1}.$ The second inequality follows from $Q$ being monotone and $g_{2}$ being the greatest selection. With an identical argument, one can prove for the case where the greatest selection from $G$ is to be increasing in $p,$ and therefore, $T$ indexed by $p,T_{p}$ is Type I (Type II) monotone in $\mathcal{M}_{1}(\mathbb{S}\times\mathbb{X}),$ with respect to $\succsim_{st}.$
@@
We are now ready to prove Theorems 1 and 2. Throughout the proof, primitive $p$ is restricted to the set $P\equiv\{p_{1},p_{2}\}$ ordered by $p_{2}\succsim p_{1}.$
@@
(ii) Proofs for Theorems 1 and 2: The first step involves showing for a given model distribution $\mu\in\mathcal{M}_{1}(\Theta),$ the set of Berk-Nash equilibrium $m$ induced by the optimal policy correspondence $G$ will be Type I (Type II) increasing in the primitives $p\in P.$ To this end, we first show that under a positive shock, the fixed point correspondence $T$ is Type I (Type II) increasing in the primitives $p.$ The fixed points of this correspondence are the Berk-Nash equilibrium $m$ for a given model distribution $\mu,$ and finally using a result in Acemoglu and Jensen (2015), we show that the set of Berk-Nash equilibrium for a given model distribution is Type I (Type II) increasing in the primitives.
@@
Because of Lemma 1, the stationary optimal policy correspondence $G$ will have a least and a greatest selection that will be increasing in $s$ and therefore, by Lemma 2, for a given model distribution $\mu,$ and primitive $p,$ $T_{\mu,p}:\mathcal{M}_{1}\left(\mathbb{S}\times\mathbb{X}\right)\rightarrow 2^{\mathcal{M}_{1}\left(\mathbb{S}\times\mathbb{X}\right)}$ induced by the stationary optimal policy correspondence, $G,$ defined in (8) is Type I (Type II) monotone with respect to $\succsim_{st}$. Now from a routine modification of Theorem B3 in Acemoglu and Jensen (2015), [^52] Theorem B3 in Acemoglu and Jensen (2015) builts on the existence theorem in Smithson (1971). It states the following: Assume that the equilibrium mapping $T$ is either Type I or type II monotone. In addition, assume that the set of measures on state and actions has an infimum. Then $T$ has fixed point. In addition, the fixed-point correspondence is upper hemicontinuous if $T$ is upper hemicontinuous. the set of fixed points, $F:\mu\times P\rightarrow 2^{\mathcal{M}_{1}\left(\mathbb{S}\times\mathbb{X}\right)}$, given by $F\left(\mu,p\right)=\left\{m\in\mathcal{M}_{1}(\mathbb{S}\times\mathbb{X}):m\in T_{\mu,p}m\right\}$ is non-empty valued and upper hemicontinuous. The proof of upper-hemicontinuity of $T_{\mu,p}$ follows for the finite case follows from Claim B (Page 744) in EP, while for the infinite case, it follows by invoking the conditions in Definition 3 on a regular SMDP, as proved in Anderson, Duanmu, Ghosh, and Khan (2023). From Lemma 2, $T_{\mu,p}$ is Type I (Type II) monotone in $p$, and therefore, from the monotonicity theorem of Acemoglu and Jensen (2015) in the main text, the set of fixed points $F$ will be non-empty and Type I (Type II) monotone in $p.$ Hence, the set of Berk-Nash equilibrium $m$ induced by the optimal policy correspondence $G$ will be Type I (Type II) increasing in the primitives $p$. This completes the first step.
@@
The second step constructs a mapping $\hat{\theta}$ that for each given model distribution $\mu$ and primitive $p\in P,$ yields a set of best-fit model distributions,
@@
A model distribution $\mu^{*}$ is an equilibrium belief if and only if $\mu^{*}\in\hat{\theta}(\mu^{*},p)$. By assumption 3 of point identification and therefore, by uniqueness, for each $m,$ there will be a non-empty unique $\theta(m).$ That is, one has a Dirac measure on $\theta(m).$ From Berge’s Maximum Theorem, $\theta$ is continuous [^53] The function $\theta_{Q}(m)$ is continuous if it is continuous in the weak*topology on its domain. and given the upper hemicontinuity of $F\left(\mu,p\right)$, $\hat{\theta}$ will be upper hemi-continuous. The rest of this step follows Acemoglu and Jensen (2015). For a fixed $\mu,$ and given $F(\mu,\cdot)$ is Type I and Type II monotone in $p,$ and increasing $\theta(m)$ under Assumption 4, one can use their Theorem 4 (pp. 601) to conclude that the least and greatest selections from $\hat{\theta}(\mu,\cdot)$ will be increasing in $p$ holding $\mu$ fixed. [^54] Theorem 4 in Acemoglu and Jensen (2015) guarantees least and greatest selections from a fixed point correspondence.

### v2

@@ 5 Analysis of Examples
In this section, we show the applicability of our results to Examples 2 and 3. Example 2 has an explicit analytical solution and is more of an illustrative example of the structure employed in this paper. However, Example 3, like most problems in dynamic programming, does not have a closed-form solution and, therefore, is most amenable to analysis with our framework.
@@
We now verify the prediction of our results vis-à-vis the analytical solution. A fall in the cost of effort $c$ is a positive shock, and therefore, leads to a higher inference of ability, $\theta^{*}=1-c,$ in the Berk-Nash equilibrium, as predicted in Theorem 2. Similarly, a fall in the cost of the effort leads to an increase in the probability of success at the steady state $m_{s}^{*}(1);$ since an increase in the model parameter is a positive shock. Here, the stationary distribution $m^{*}_{s}$ is increasing in the usual (first) stochastic order dominance in the cost of effort.
@@ 6 Welfare Comparisons
We now turn towards comparing the objective welfare of an agent facing a regular SMDP, under correctly specified and misspecified Bayesian learning and characterize an upper bound on the difference between the two instances, the costs of misspecification, in terms of the primitives of the environment. For the purposes of this section, we shall assume that the utility function is bounded and strictly concave. [^52] Another simplification we use here is working with payoff functions that have the current state and action as arguments. Further, we shall assume that $u:\mathbb{S}\times\mathbb{X}\rightarrow\mathbb{R}$ is continuously differentiable in actions.
Let $g(s,\bar{\theta}),$ parameterized by a parameter $\bar{\theta},$ denote an optimal policy function. Suppose it is well-defined under both correctly specified and misspecified environments, where $\bar{\theta}$ take values $\theta^{*}$ and $\theta_{*}$ under correctly specified and misspecified learning, respectively. We are interested in comparing the welfare of a correctly specified agent with parameter $\theta^{*}$ to the welfare of an agent who settles at the parameter $\theta_{*}$ asymptotically, under misspecified Bayesian learning as in Equation (1). Then, the agent’s welfare under the parameter $\bar{\theta},$ is the objective ex-ante expected discounted payoff, $W(s,\bar{\theta})$ of choosing the optimal action, $g(s,\bar{\theta}),$ and is given as,
@@
We interpret the difference $\gamma$ between the two policy functions in the sup-norm as the approximation error in the space of policy functions. Then, for a given approximation error $\gamma,$ Equation (17) supplies an upper bound on the welfare comparison in terms of the primitives, namely the discount factor, and the absolute bounds on the utility, the marginal utility function, and the KL divergence. Notice that if $\gamma=0,$ that is, the policy functions under the two instances of learning are identical, then the two welfare quantities are equal, by construction. The upper bound in Equation (17) gives intuitive comparative statics, vis-à-vis the model primitives. For instance, the discrepancy in welfare is larger the greater the approximation error, $\gamma.$ Similarly, for a given positive approximation error, the discrepancy increases in the discount factor. That is, as the agent gets increasingly patient–reflected in a higher discount factor $\beta$–his approximation error accumulates over the horizon, leading to a greater discrepancy in welfare. Furthermore, higher bounds on the utility function and its corresponding marginal utility function lead to a greater discrepancy in the welfare. The marginal utility function plays a role in the discrepancy by attaching itself to the approximation error. Intuitively, getting the policy wrong is more pronounced if it matters more at the margins. Notice that for the upper bound on the KL divergence, we only compare the implied distributions by the true transition function $Q$ under the two policy functions. This is because our interest lies in comparing the objective welfare under $Q$.
It is worth mentioning here that bound in Equation (17) is not specific to a misspecified MDP, rather, it holds for any MDP environment with a endogenous state evolution process, when plugged with alternate policy functions. In this sense, it is related to Lemma 3.1 in Santos (2000) which does a similar exercise in bounding welfare discrepancies. However, the state process in there is exogeneous. Furthermore, Santos (2000) uses Euler residuals to provide upper bounds on the approximation error in terms of the primitives for both the policy and the value functions, which is useful for computing numerical approximations. Often, MDP environments do not have closed form solutions, thus making economists rely on numerical techniques to solve for them computationally. We hope that Theorem 5 is instructive in this regard and helps initiate conversations on the numerical approximation of Berk-Nash equilibria.
@@ 7 Concluding Remarks and Extensions
Models, by their very nature, offer simplified abstractions of reality, inevitably omitting certain nuances. This paper outlines conditions under which an important qualitative property, the comparative statics of decision-making and the corresponding inference, is preserved under model misspecification. The main contribution of this paper lies in establishing monotone comparative statics results for misspecified dynamic optimization problems, by a novel application of techniques in the fixed points literature (Smithson (1971)) and first used in the context of large dynamic economies (Acemoglu and Jensen (2015)). We also illustrate the utility of these results for general interest environments. Further, we provide an upper bound on the welfare comparison between correct and misspecified learning, in terms of the primitives. We conclude by outlining several promising directions for extending our findings to more broader environments.
@@
Multi-dimensional parameter spaces. Since most of the current applications in the literature predominantly involve one-dimensional models, [^54] This feature of the current literature is also noted in Esponda, Pouzo, and Yamamoto (2021). the results presented here focus on the model (parameter) set as a compact subset of the real line. However, they could be extended to multi-dimensional parameter spaces. This would involve a suitable rehabilitation of Assumptions 3 and 4 for lattice and non-lattice multi-dimensional parameter spaces. Although one limitation of these results using the current techniques is that the comparative statics results imply all inferred parameters in equilibrium must be monotonic together, which may not be ideal for specific economic applications.
@@
Other forms of updating. While our results primarily address the asymptotic behavior of Bayesian learners, it’s important to note that other inference processes, such as Maximum Likelihood Estimation (MLE) and moment-based learning, can lead to the same asymptotic inferences of best-fit models. Csaba and Szoke (2023) provide an illustrative example that demonstrates the identical nature of inference between Bayesian inference and methods based on likelihoods.
@@
Static environments and dynamic concerns. Our results are tailored for dynamic MDP environments but they potentially are also applicable, subject to modifications, to static environments, such as the one in Esponda and Pouzo (2016) for static games. This also spills over to settings where agents are endogenously concerned about their misspecified models, such as those in Lanzani (2025). [^55] Corollary 1 in Lanzani (2025) presents an interesting comparative statics application to monetary policy cycles; see it for more details.
@@
Monotone comparative dynamics. The scope of our paper is limited to analyzing the comparative statics properties of the equilibrium objects in the steady state. An interesting question arises regarding how the dynamics of the misspecified learning process might react to changes in its primitives. Specifically, we do not know how the variations in the frequency of state-action pairs, as well as the sequence of posteriors as described in Equation (1), would be influenced because of changes in the economic primitives. There is added complexity in the dynamics due to potential complementarities between current actions and the inference process via the role of experimentation. In this regard, the tools developed in Balbus, Dziewulski, Reffett, and Woźny (2022) for monotone comparative dynamics for stochastic games could be useful for further exploration.
@@
Misspecification in large economies. Our results are also potentially applicable to misspecified dynamic economies with a continuum of agents, such as the one conceived in Molavi (2019). Given that the techniques we rely on (Acemoglu and Jensen (2015)) were developed in the context of economies with a continuum of agents, the results in this paper should potentially be applicable on equilibrium concepts that Molavi (2019) develops for the boundedly rational agents working with misspecified models in macroeconomic environments.
@@ Appendix B Proofs
The proofs for Theorems 1 and 2 are structured in three steps, with the proof technique similar to that in Acemoglu and Jensen (2015). For Theorem 2, the first step involves showing that for any fixed model distribution $\mu\in\mathcal{M}_{1}(\Theta)$, the set of stationary distributions on states and actions, $m^{*}$, induced by the optimal policy correspondence $G,$ will be Type I (Type II) monotonic in the primitives $p$. The second step involves constructing a mapping $\hat{\theta}$ that for each given $\mu$ and $p$ yields a set of model distributions, $\mu^{\prime}s$. It is the fixed points of this map that are the equilibrium model distributions $\mu^{*},$ given $p$. Finally, the third step involves the least and greatest selections from this map will be increasing in $p.$ This in turn leads us to give a new existence proof of Theorem 1 that relies on the monotonicity and identification properties of the equilibrium map, $T$. The rest of the proofs (Theorems 3 and 4) shall follow an analogous structure. Proof of Theorem 5 is based on the Taylor expansion of the utility function and on an application of a new result proven by Canonne (2022) on entropy bounds.
The equilibrium mapping $T$ associated with the Berk-Nash equilibrium (Definition 5) is a set-valued function on the product space of probability measure on states and actions, and parameter space, $T:W\rightarrow 2^{W},$ where $W=\mathcal{M}_{1}(\mathbb{S}\times\mathbb{X})\times\mathcal{M}_{1}(\Theta)$ such that $\mathcal{T}(m,\mu)=\mathcal{M}(m,\mu)\times\mathcal{M}\left(\Theta_{Q}(m)\right)$, where
@@
for any $\mu\in\mathcal{M}_{1}(\Theta),\mathcal{F}(\mu)$ is the set of all $m^{\prime}$ that satisfy the conditions of optimality and stationarity as defined in Definition 5.
@@
Let the regular SMDP satisfy assumptions 1 & 2. Then for any given primitive $p,$ the optimal policy correspondence $G:\mathbb{S}\times P\rightarrow 2^{\mathbb{X}}$ is increasing in the state $s$ in the strong set order. In particular, it has a least and a greatest selection, and they are increasing in state $s$ [^57] We suppress the dependence on $\mathcal{M}_{1}(\Theta)$ for notational convenience. Theorem 3.9.2 in Topkis (1998) establishes conditions under which finite and infinite period discounted MDPs have increasing policy selection.
@@
$G(s,\mu,p)=\operatornamewithlimits{argmax}\limits_{x\in\mathbb{X}}\displaystyle\ \int_{S}\{u(s,x,s^{\prime})+\delta\ V(s^{\prime},\mu,p)\}\bar{Q_{\mu}}(\mathrm{d}s^{\prime}|s,x),$ (19)
where $\bar{Q}_{\mu}=\displaystyle\int_{\Theta}Q_{\theta}\mu(d\theta).$ Given Assumptions 1 and 2, and by the application of Theorem 3.9.2 (page 165, Topkis (1998)), the expression on the right-hand side is supermodular in $(s,x)$. Therefore, the optimal policy correspondence will be increasing in $s$ in the strong-set order. Further, given the optimal correspondence, the greatest and least selection exist and are increasing in the state.
@@
We only prove the above statement for the greatest selection in the Type I case. [^58] The rest of the cases involving least selection and Type II monotonicity follow analogously. Consider probability measures, $\nu_{1},\nu_{2}\in\mathcal{M}_{1}(\mathbb{S}\times\mathbb{X})$ such that $\nu_{2}\succsim_{st}\nu_{1}.$ To prove that the fixed point correspondence $T$ is Type I monotone, we need to show that for any $\lambda_{1}\in$ $T\nu_{1}$, there exists a $\lambda_{2}\in T\nu_{2}$ such that $\lambda_{2}\succsim_{st}\lambda_{1}.$ That is, if $\lambda_{1}\in T\nu_{1},$ then there exists a measurable selection $g_{1}:\mathbb{S}\rightarrow\mathbb{X},$ such that, for all $A\times B\in\mathcal{B}(\mathbb{S}\times\mathbb{X})$,
@@
The first inequality follows from $\nu_{2}\succsim_{st}\nu_{1}.$ The second inequality follows from $Q$ being monotone and $g_{2}$ being the greatest selection. With an identical argument, one can prove for the case where the greatest selection from $G$ is to be increasing in primitive $p,$ and therefore, $T$ indexed by $p,T_{p}$ is Type I (Type II) monotone in $\mathcal{M}_{1}(\mathbb{S}\times\mathbb{X}),$ with respect to $\succsim_{st}.$
@@
We are now ready to prove Theorems 1 and 2. Throughout the proof, and without loss of generality, primitive $p$ is restricted to the set $P\equiv\{p_{1},p_{2}\}$ ordered by $p_{2}\succsim p_{1},$ where the ordering depends on the primitive being considered.
@@
(ii) Proofs for Theorems 1 and 2: The first step involves showing for a given model distribution $\mu\in\mathcal{M}_{1}(\Theta),$ the set of Berk-Nash equilibriua $m$ induced by the optimal policy correspondence $G$ will be Type I (Type II) increasing in the primitives $p\in P.$ To this end, we first show that under a positive shock, the fixed point correspondence $T$ is Type I (Type II) increasing in the primitives $p.$ The fixed points of this correspondence are the Berk-Nash equilibrium $m$ for a given model distribution $\mu,$ and finally using a result in Acemoglu and Jensen (2015), we show that the set of Berk-Nash equilibriua for a given model distribution is Type I (Type II) increasing in the primitives.
@@
Because of Lemma 1, the stationary optimal policy correspondence $G$ will have a least and a greatest selection that will be increasing in $s$ and therefore, by Lemma 2, for a given model distribution $\mu,$ and primitive $p,$ $T_{\mu,p}:\mathcal{M}_{1}\left(\mathbb{S}\times\mathbb{X}\right)\rightarrow 2^{\mathcal{M}_{1}\left(\mathbb{S}\times\mathbb{X}\right)}$ induced by the stationary optimal policy correspondence, $G,$ defined in (8) is Type I (Type II) monotone with respect to $\succsim_{st}$. Now from a routine modification of Theorem B3 in Acemoglu and Jensen (2015), [^59] Theorem B3 in Acemoglu and Jensen (2015) builts on the existence theorem in Smithson (1971). It states the following: Assume that the equilibrium mapping $T$ is either Type I or type II monotone. In addition, assume that the set of measures on state and actions has an infimum. Then $T$ has fixed point. In addition, the fixed-point correspondence is upper hemicontinuous if $T$ is upper hemicontinuous. the set of fixed points, $F:\mu\times P\rightarrow 2^{\mathcal{M}_{1}\left(\mathbb{S}\times\mathbb{X}\right)}$, given by $F\left(\mu,p\right)=\left\{m\in\mathcal{M}_{1}(\mathbb{S}\times\mathbb{X}):m\in T_{\mu,p}m\right\}$ is non-empty valued and upper hemicontinuous. The proof of upper-hemicontinuity of $T_{\mu,p}$ follows for the finite case follows from Claim B (page 744) in EP, while for the infinite case, specifically compact Euclidean spaces here, it follows by invoking the conditions in Definition 3 on a regular SMDP, as proved in Anderson, Duanmu, Ghosh, and Khan (2024). From Lemma 2, $T_{\mu,p}$ is Type I (Type II) monotone in $p$, and therefore, from the monotonicity theorem of Acemoglu and Jensen (2015) in the main text (page 15), the set of fixed points $F$ will be non-empty and Type I (Type II) monotone in $p.$ Hence, the set of Berk-Nash equilibriua $m$ induced by the optimal policy correspondence $G$ will be Type I (Type II) increasing in the primitives $p$. This completes the first step.
@@
The second step constructs a mapping $\hat{\theta}:\mathcal{M}_{1}(\Theta)\times P\rightarrow 2^{\mathcal{M}_{1}(\Theta)}$ that for each given model distribution $\mu$ and primitive $p\in P,$ yields a set of best-fit model distributions,
@@
A model distribution $\mu^{*}$ is an equilibrium belief if and only if $\mu^{*}\in\hat{\theta}(\mu^{*},p)$. By assumption 3 of point identification and therefore, by uniqueness, for each $m,$ there will be a non-empty unique $\theta(m).$ That is, one has a Dirac measure on $\theta(m).$ From Berge’s Maximum Theorem, $\theta$ is continuous [^60] The function $\theta_{Q}(m)$ is continuous if it is continuous in the weak*topology on its domain. and given the upper hemicontinuity of $F\left(\mu,p\right)$, $\hat{\theta}$ will be upper hemi-continuous. The rest of this step follows Acemoglu and Jensen (2015). For a fixed $\mu,$ and given $F(\mu,\cdot)$ is Type I and Type II monotone in $p,$ and increasing $\theta(m)$ under Assumption 4, one can use their Theorem 4 (pp. 601) to conclude that the least and greatest selections from $\hat{\theta}(\mu,\cdot)$ will be increasing in $p$ holding $\mu$ fixed. [^61] Theorem 4 in Acemoglu and Jensen (2015) guarantees least and greatest selections from a fixed point correspondence. It may be possible to dispense with identification (Assumption 3) if we were to use more general divergence measures apart from the Kullback-Leibler divergence. However, given the criticality of the KL divergence, uniqueness is a feature we require. It is also something that the current literature has emphasized in terms of applications; see Esponda and Pouzo (2021); Esponda, Pouzo, and Yamamoto (2021).